MW SYSTEMATIC STUDY OF ALKALOIDS: THE DISTORTED TROPANE OF SCOPOLINE

\underline{PATRICIA ECIJA}, EMILIO J. COCINERO, FRANCISCO J. BASTERRETXEA, JOSÉ A. FER-NANDEZ, FERNANDO CASTANO, Departamento de Química Física, Facultad de Ciencia y Tecnología, Universidad del País Vasco (UPV/EHU), Ap.644, E-48940, Bilbao, Spain; ALBERTO LESARRI, Departamento de Química Física y Química Inorgánica, Facultad de Ciencias, Universidad de Valladolid, E-47011 Spain.

Tropane alkaloids have diverse pharmacological uses and are well-known for their neurostimulant activity. Previous structure-activity-relationship established correlations between bioactivity and several aspects of ligand conformation and stereochemistry, including delicate intramolecular effects like nitrogen inversion^{a}. We have initiated a series of structural studies on tropane alkaloids^{b}, aimed to discerning their intrinsic stereochemical properties using rotational spectroscopy in supersonic jets^{c}. Here we extend these studies to the epoxytropanes, initially motivated to interrogate the influence of the epoxy group on nitrogen inversion and ring conformation. The rotational spectrum evidences a single structure in the gas phase, providing a first description of the (three ring) structurally-distorted tropane in scopoline. The determined rotational parameters of scopoline reveal the structural consequences of the intramolecular cyclation of scopine, which breaks the original epoxy group and creates a new ether bridge and a 7β-hydroxytropane configuration. The hydroxyl group further stabilizes the molecule by an O−H ⋯ N intramolecular hydrogen bond, which, in turn, forces the N-methyl group to the less stable axial form^{b}. The experimental work was supported by ab initio and DFT calculations.

a i) S.Singh, Chem. Rev. 100, 925 (2000); ii) A. Krunic, D. Pan, W.J. Dunn III, S.V.S. Miariappan, Bioorg. Med. Chem. 17, 811 (2009). b E.J. Cocinero, A. Lesarri, P. Écija, J.-U. Grabow, J.A. Fernández, F. Castaño, Phys. Chem. Chem. Phys. 12, 6076 (2010). c E.J. Cocinero, A. Lesarri, P. Écija, J.-U. Grabow, J.A. Fernández, F. Castaño, Phys. Chem. Chem. Phys. 12, 12486